A ship off-loaded three different types of high-attitude, high-performance jet engines namely Turbofan, Pulse Jet and Scram Jet. It was found that Turbofan engine weighed 7kgs more than Scramjet engine and Scramjet engine weighed 20kg less than Pulse jet. If the total weight of the three engines is 1248kg, find the weight of each type of engine.
Answer by ptaylor(1894)

Let x=weight of Pulse Jet
Then (x-20)=weight of Scramjet
And (x-20)+7=(x-13)=weight of Turbofan
Now we are told that:
x+(x-20)+(x-13)=1248 or
x+x-20+x-13=1248 simplify
3x-33=1248 add 33 to both sides
3x-33+33=1248+33 collect like terms
3x=1281 divide both sides by 3
x=427 kg--------------------weight of Pulse Jet
x-20=427-20=407 kg-------------------------weight of Scramjet
x-13=427-13=414 kg------------------------------weight of Turbofan
CK
427+407+414=1248
1248=1248

Hope this helps---ptaylor